# High-dimensional analysis of ridge regression for non-identically distributed data with a variance profile

### v2

@@ 5 Conclusion
$\hat{R}(\hat{\theta}_{\lambda},\theta_{\ast})$ $=$ $\mathbb{E}[\lambda^{2}\theta_{\ast}^{t}Q_{p}(-\lambda)\widetilde{S}_{p}Q_{p}(-\lambda)\theta_{\ast}+\frac{\varepsilon_{n}^{t}X_{n}}{n}Q_{p}(-\lambda)\widetilde{S}_{p}Q_{p}(-\lambda)\frac{X_{n}^{t}\varepsilon_{n}}{n}$
$-2\lambda\theta_{\ast}^{t}Q_{p}(-\lambda)\widetilde{S}_{p}Q_{p}(-\lambda)\frac{X_{n}^{t}\varepsilon_{n}}{n}|X_{n}]$
$=$ $\lambda^{2}\mathrm{Tr}[Q_{p}(-\lambda)\widetilde{S}_{p}Q_{p}(-\lambda)\mathbb{E}[\theta_{\ast}\theta_{\ast}^{t}]]$
$+\frac{1}{n^{2}}\mathrm{Tr}[\widetilde{S}_{p}Q_{p}(-\lambda)X_{n}^{t}\mathbb{E}[\varepsilon_{n}\varepsilon_{n}^{t}]X_{n}Q_{p}(-\lambda)]$
@@
Note that $\widehat{\Sigma}_{n}$ and $Q_{p}(-\lambda)$ are commuting, one has that $Q_{p}^{2}(-\lambda)=Q_{p}^{\prime}(-\lambda)$, and moreover, $\widehat{\Sigma}_{n}Q_{p}^{\prime}(-\lambda)=Q_{p}(-\lambda)-\lambda Q_{p}^{\prime}(\lambda)$. Then we get the following Bias-Variance decomposition of the predictive risk
$\hat{R}(\hat{\theta}_{\lambda},\theta_{\ast})$ $=$ $\mathrm{Bias}(\hat{\theta}_{\lambda})+\mathrm{Var}(\hat{\theta}_{\lambda}),$
where
$\mathop{\rm Bias}\nolimits(\hat{\theta}_{\lambda})=\frac{\lambda^{2}\alpha^{2}}{p}\mathrm{Tr}[Q_{p}(-\lambda)\widetilde{S}_{p}Q_{p}(-\lambda)]=\frac{\lambda^{2}\alpha^{2}}{p}\mathrm{Tr}[\widetilde{S}_{p}Q_{p}^{\prime}(-\lambda)]$
and
$\mathop{\rm Var}(\hat{\theta}_{\lambda})=\frac{\sigma^{2}}{n}\operatorname{Tr}[\widetilde{S}_{p}\widehat{\Sigma}_{n}Q_{p}^{\prime}(-\lambda)]=\frac{\sigma^{2}}{n}\mathrm{Tr}[\widetilde{S}_{p}(Q_{p}(-\lambda)-\lambda Q_{p}^{\prime}(-\lambda))].$
@@
$\hat{r}_{\lambda}(X_{n})=\sigma^{2}+\frac{\sigma^{2}}{n}\mathrm{Tr}[\widetilde{S}_{p}Q_{p}(-\lambda)]+\lambda\left(\frac{\lambda\alpha^{2}}{p}-\frac{\sigma^{2}}{n}\right)\mathrm{Tr}[\widetilde{S}_{p}Q_{p}^{\prime}(-\lambda)],$
@@
A.2 Deterministic equivalents of operator-valued Stieltjes transforms
One can remark that the deterministic equivalent given in Equation (3.4) is obtained by replacing $\Delta[Q_{p}(-\lambda)]$ with $T_{p}(-\lambda)$ in Equation (3.3). As argued in Section 3.2, even though Proposition 2.1 states that $\frac{1}{n}\operatorname{Tr}[Q_{p}(-\lambda)]\sim\frac{1}{n}\operatorname{Tr}[T_{p}(-\lambda)]$, this equivalence is not sufficient to prove Theorem 3.1. We need to find a deterministic matrix that is a good approximation of the diagonal of the resolvent $\Delta[Q_{p}(z)]$ in a high-dimensional sense. To this end, we first define the following equivalence relation in order to specify an appropriate notion of deterministic equivalent for the matrix $\Delta[Q_{p}(-\lambda)]$.
Definition A.1.
for all family of deterministic matrices $(U_{p})_{p\geq 1}$ satisfying :
$\forall p\geq 1,\ \ \ U_{p}\in\mathbb{R}^{p\times p}$ $\mathrm{and}$ $\exists K>0,\ \sup_{p\geq 1}\max_{i\in[p]}|U_{p}^{(i)}|\leq K,$ (A.3)
where $U_{p}^{(i)}$ denotes the i-th diagonal entry of $U_{p}$.
Then, using Definition A.1, we prove that $T_{p}(z)$ is a relevant deterministic equivalent of $Q_{p}(z)$ in the following sense.
Theorem A.1.
Let us define the following families of matrices $\Delta[\mathbf{Q}(z)]=(\Delta[Q_{p}(z)])_{p\geq 1}$, $\Delta[\mathbf{Q^{\prime}}(z)]=(\Delta[Q^{\prime}_{p}(z)])_{p\geq 1}$, $\mathbf{T}(z)=(T_{p}(z))_{p\geq 1}$ and $\mathbf{T^{\prime}}(z)=(T^{\prime}_{p}(z))_{p\geq 1}$. Then, the following equivalences hold:
Theorem A.1 is inspired by results from [HLN07], and it constitutes a key element in the proof of Theorem 3.1. In order to prove Theorem A.1, we introduce the diagonal matrices $R_{p}(z)=\underset{1\leq j\leq p}{\mathop{\rm diag}\nolimits}(R^{(j)}_{p}(z))$ defined by
@@ A.2 Deterministic equivalents of operator-valued Stieltjes transforms
with $\widetilde{Q}_{n}(z)=\Big{(}\widetilde{\Sigma}_{n}-zI_{n}\Big{)}^{-1}$ is the resolvent matrix of $\widetilde{\Sigma}_{n}=\frac{1}{n}X_{n}X_{n}^{t}$.
@@
Let us define $\Delta[\mathbf{Q}(z)]$ as in Theorem A.1 and denote by $\mathbf{R}(z)$ the family of matrices $(R_{p}(z))_{p\geq 1}$. Then, the following equivalences hold:
$\Delta[\mathbf{Q}(z)]\sim\mathbf{R}(z),$
for all $z\in\mathbb{C}^{+}=\{z\in\mathbb{C},\Im(z)>0\}$.
@@
This proof is based on [HLN07][Lemma 6.1] which asserts that there exists a constant $K_{1}>0$ such that $\mathbb{E}\left[\left|\frac{1}{p}\operatorname{Tr}[(Q_{p}(z)-R_{p}(z))U_{p}]\right|^{2+\delta/2}\right]\leq\frac{K_{1}}{p^{1+\delta/4}}$, where $(U_{p})_{p\geq 1}$ denotes a family of deterministic matrices satisfying (A.3). We deduce from this inequality that
@@
The serie being finite, we obtain that $\mathbb{P}\left[\sum\limits_{p\geq 1}\left|\frac{1}{p}\operatorname{Tr}[(Q_{p}(z)-R_{p}(z))U_{p}]\right|^{2+\varepsilon/2}<+\infty\right]=1$, which finally gives us :
$\mathbb{P}\left[\frac{1}{p}\operatorname{Tr}[(Q_{p}(z)-R_{p}(z))U_{p}]\xrightarrow[n\rightarrow+\infty]{}0\right]\geq\mathbb{P}\left[\sum_{p\geq 1}\left|\frac{1}{p}\operatorname{Tr}[(Q_{p}(z)-R_{p}(z))U_{p}]\right|^{2+\varepsilon/2}<+\infty\right]=1.$
@@
Then, Lemma A.1 needs to be combined with the following one in order to prove Theorem A.1.
@@
Let us define $\mathbf{T}(z)$ and $\mathbf{R}(z)$ as in Theorem A.1 and Lemma A.1. Then, the following equivalences hold:
$\mathbf{R}(z)\sim\mathbf{T}(z),$
for all $z\in\mathcal{D}=\{z\in\mathbb{C}^{+},\frac{|z|}{|\Im(z)|}<2,|\Im(z)|>4d\gamma_{max}^{2}\}$.
@@
Since $T_{p}(z)$ and $R_{p}(z)$ are diagonal matrices, the following equations hold true
$\mathrm{Tr}[(R_{p}(z)-T_{p}(z))U_{p}]=\sum_{i=1}^{p}(R_{p}^{(i)}(z)-T_{p}^{(i)}(z))U_{p}^{(i)},$
where $(U_{p})_{p\geq 1}$ denotes a family of deterministic matrices satisfying (A.3), and
@@
We know from [HLN07][Proposition 5.1], that $|R_{p}^{(i)}(z)T_{p}^{(i)}(z)|\leq\frac{1}{|\Im(z)|^{2}}$. Moreover [HLN07][Equation (6.15)] states that
@@
Therefore, we obtain from these inequalities and (A.3) that
$\mathbb{E}\left[\sum_{p\geq 1}\left|\frac{1}{p}\operatorname{Tr}[(T_{p}(z)-R_{p}(z))U_{p}]\right|^{2+\delta/2}\right]$$\leq$ $\sum_{p\geq 1}\frac{1}{p}\sum_{i=1}^{p}|U_{p}^{(i)}|^{2+\delta/2}\mathbb{E}\left[|T_{p}^{(i)}(z)-R_{p}^{(i)}(z)|^{2+\delta/2}\right]$
$\leq$ $\sum_{p\geq 1}\frac{K^{2+\delta/2}K_{2}}{|\Im(z)|^{2}p^{1+\delta/4}}$
$\leq$ $\sum_{p\geq 1}\frac{\widetilde{K}}{p^{1+\delta/4}}$
$<$ ++$\infty$,+ ∞ ,
@@
Proof of Theorem A.1.
Since the equivalence relation introduced in Definition A.1 is transitive, we deduce from Lemma A.1 and Lemma A.2 that
$\frac{1}{p}\operatorname{Tr}[Q_{p}(z)U_{p}]\sim\frac{1}{p}\operatorname{Tr}[T_{p}(z)U_{p}],$ (A.4)
for $z\in\mathcal{D}$. It remains to prove that this equivalence is true for $z\in\mathbb{C}\setminus\mathbb{R}^{+}$. To this end, let us define the function
$f_{p}(z)=\frac{1}{p}\operatorname{Tr}[Q_{p}(z)U_{p}]-\frac{1}{p}\operatorname{Tr}[T_{p}(z)U_{p}]$
for $z\in\mathbb{C}\setminus\mathbb{R}^{+}$. Proving that (A.4) holds for $z\in\mathbb{C}\setminus\mathbb{R}^{+}$ is equivalent to prove that $f_{p}(z)$ converges uniformly to zero. This function is analytic on $z\in\mathbb{C}\setminus\mathbb{R}^{+}$ since $Q_{p}(z)$ and $T_{p}(z)$ are analytic [HLN07][Proposition 5.1]. For $1\leq i\leq p$, $T_{p}^{(i)}(z)$ is a Stieltjes transforms [HLN07][Theorem 2.4], then we can state that $|T_{p}^{(i)}(z)|\leq\frac{1}{\mathrm{dist}(z,\mathbb{R}^{+})}$ with $z\in\mathbb{C}\setminus\mathbb{R}^{+}$, according to [HLN07][Proposition 2.2]. Moreover, we deduce from [HLN07][Proposition 2.3] and [HS12][Theorem 5.8] that $|Q_{p}^{(i)}(z)|\leq\parallel Q_{p}(z)\parallel_{sp}\leq\frac{1}{\mathrm{dist}(z,\mathbb{R}^{+})}$, where .$\parallel.\parallel_{sp}$ is the spectral norm, $Q_{p}^{(i)}(z)$ denotes the i-th diagonal entry of $Q_{p}(z)$ and $z\in\mathbb{C}\setminus\mathbb{R}^{+}$. Thus, using (A.3), we have proved that
@@
Hence, for each compact subset $C\subset\mathbb{C}\setminus\mathbb{R}^{+}$, $f_{p}$ is uniformly bounded on $C$,
$|f_{n}(z)|\leq\frac{2K}{\delta_{C}},$
where $\delta_{C}$ is the distance between $C$ and $\mathbb{R}^{+}$. Then, by the normal family Theorem [Rud87][Theorem 14.6] there exists a sub-sequence $f_{p_{k}}$ which uniformly converges to $f^{\ast}$ that is an analytical function on $\mathbb{C}\setminus\mathbb{R}^{+}$. Let $(z_{k})_{k\in\mathbb{N}}$ be a sequence with an accumulation point in $\mathcal{D}$. Then, for each $k$, $f_{p}(z_{k})\rightarrow 0$ with probability one as $p$ tends to +$\infty$+ ∞. This implies that $f^{\ast}(z_{k})=0$ for each $k$. We finally obtain that $f^{\ast}$ is identically zero on $\mathbb{C}\setminus\mathbb{R}^{+}$. Therefore $f_{p}$ converges uniformly to zero which proves that (A.4) holds for $z\in\mathbb{C}\setminus\mathbb{R}^{+}$.
Now, it remains to prove that
$\frac{1}{p}\operatorname{Tr}[Q_{p}^{\prime}(z)U_{p}]\sim\frac{1}{p}\operatorname{Tr}[T^{\prime}(z)U_{p}].$ (A.6)
The functions $Q_{p}(z)$ and $T_{p}(z)$ being analytic on $\mathbb{C}\setminus\mathbb{R}^{+}$ [HS12][Theorem 1.2],[HLN07][Proposition 5.1], the Cauchy integral formula yields that
@@
Since we have proved that (A.4) holds on $\mathbb{C}\setminus\mathbb{R}^{+}$, we obtain that for all $w\in\rho$
$\lim_{n\to\infty,\;p/n\to c}\frac{1}{p}\mathrm{Tr}[(Q_{p}(w)-T_{p}(w))U_{p}]=0.$
Moreover, for $w\in\mathbb{C}\setminus\mathbb{R}^{+}$, we have from (A.5) that
@@
As $\frac{2K}{\mathrm{dist}(\rho,\mathbb{R}^{+})|w-z|^{2}}$ is integrable on $\rho$, using the dominated convergence theorem, we obtain that (A.6) holds for $z\in\mathbb{C}\setminus\mathbb{R}^{+}$ which completes the proof of Theorem A.1. ∎

### v3

@@ 5 Conclusion
$\hat{R}(\hat{\theta}_{\lambda},\beta_{\ast})$ $=$ $\mathbb{E}[\lambda^{2}\beta_{\ast}^{\top}Q_{p}(-\lambda)\widetilde{S}_{p}Q_{p}(-\lambda)\beta_{\ast}+\frac{\varepsilon_{n}^{\top}X_{n}}{n}Q_{p}(-\lambda)\widetilde{S}_{p}Q_{p}(-\lambda)\frac{X_{n}^{\top}\varepsilon_{n}}{n}$
$-2\lambda\beta_{\ast}^{\top}Q_{p}(-\lambda)\widetilde{S}_{p}Q_{p}(-\lambda)\frac{X_{n}^{\top}\varepsilon_{n}}{n}|X_{n}]$
$=$ $\lambda^{2}\mathrm{Tr}[Q_{p}(-\lambda)\widetilde{S}_{p}Q_{p}(-\lambda)\mathbb{E}[\beta_{\ast}\beta_{\ast}^{\top}]]$
$+\frac{1}{n^{2}}\mathrm{Tr}[\widetilde{S}_{p}Q_{p}(-\lambda)X_{n}^{\top}\mathbb{E}[\varepsilon_{n}\varepsilon_{n}^{\top}]X_{n}Q_{p}(-\lambda)]$
@@
Note that $\widehat{\Sigma}_{n}$ and $Q_{p}(-\lambda)$ are commuting, one has that $Q_{p}^{2}(-\lambda)=Q_{p}^{\prime}(-\lambda)$, and moreover, $\widehat{\Sigma}_{n}Q_{p}^{\prime}(-\lambda)=Q_{p}(-\lambda)-\lambda Q_{p}^{\prime}(\lambda)$. Then we get the following Bias-Variance decomposition of the predictive risk $\hat{R}(\hat{\theta}_{\lambda},\beta_{\ast})=\mathrm{Bias}(\hat{\theta}_{\lambda})+\mathrm{Var}(\hat{\theta}_{\lambda}),$ where $\mathop{\rm Bias}\nolimits(\hat{\theta}_{\lambda})=\frac{\lambda^{2}\alpha^{2}}{p}\mathrm{Tr}[Q_{p}(-\lambda)\widetilde{S}_{p}Q_{p}(-\lambda)]=\frac{\lambda^{2}\alpha^{2}}{p}\mathrm{Tr}[\widetilde{S}_{p}Q_{p}^{\prime}(-\lambda)]$
and $\mathop{\rm Var}(\hat{\theta}_{\lambda})=\frac{\sigma^{2}}{n}\operatorname{Tr}[\widetilde{S}_{p}\widehat{\Sigma}_{n}Q_{p}^{\prime}(-\lambda)]=\frac{\sigma^{2}}{n}\mathrm{Tr}[\widetilde{S}_{p}(Q_{p}(-\lambda)-\lambda Q_{p}^{\prime}(-\lambda))].$
@@
$\hat{r}_{\lambda}^{test}(X_{n})=\sigma^{2}+\frac{\sigma^{2}}{n}\mathrm{Tr}[\widetilde{S}_{p}Q_{p}(-\lambda)]+\lambda\left(\frac{\lambda\alpha^{2}}{p}-\frac{\sigma^{2}}{n}\right)\mathrm{Tr}[\widetilde{S}_{p}Q_{p}^{\prime}(-\lambda)],$
@@
A.2 Derivation of the deterministic equivalent the diagonal of the resolvent
This section is dedicated to prove Theorem 3.1 that aims to provide a deterministic equivalent of $\Delta[Q_{p}(z)]$ that is necessary to obtain asymptotic equivalents of the risks. The proof of this theorem is strongly inspired by results from [HLN07], and it constitutes a key element in the proof of Theorem 3.2. In order to prove Theorem 3.1, we introduce the diagonal matrices $R_{p}(z)=\underset{1\leq j\leq p}{\mathop{\rm diag}\nolimits}(R^{(j)}_{p}(z))$ defined by
@@ A.2 Derivation of the deterministic equivalent the diagonal of the resolvent
with $\widetilde{Q}_{n}(z)=\Big{(}\widetilde{\Sigma}_{n}-zI_{n}\Big{)}^{-1}$ is the resolvent matrix of $\widetilde{\Sigma}_{n}=\frac{1}{n}X_{n}X_{n}^{\top}$.
@@
Let us define $\Delta[\mathbf{Q}(z)]$ as in Theorem 3.1 and denote by $\mathbf{R}(z)$ the family of matrices $(R_{p}(z))_{p\geq 1}$. Then one has that $\Delta[\mathbf{Q}(z)]\sim\mathbf{R}(z),$ for all $z\in\mathbb{C}^{+}=\{z\in\mathbb{C},\Im(z)>0\}$.
@@
This proof is based on [HLN07][Lemma 6.1] which asserts that there exists a constant $K_{1}>0$ such that $\mathbb{E}\left[\left|\frac{1}{p}\operatorname{Tr}[(Q_{p}(z)-R_{p}(z))U_{p}]\right|^{2+\delta/2}\right]\leq\frac{K_{1}}{p^{1+\delta/4}}$, where $(U_{p})_{p\geq 1}$ denotes a family of deterministic matrices satisfying (3.3). We deduce from this inequality that
@@
The serie being finite, one has that
$\mathbb{P}\left[\sum\limits_{p\geq 1}\left|\frac{1}{p}\operatorname{Tr}[(Q_{p}(z)-R_{p}(z))U_{p}]\right|^{2+\varepsilon/2}<+\infty\right]=1$
, which finally gives us:
$\mathbb{P}\left[\frac{1}{p}\operatorname{Tr}[(Q_{p}(z)-R_{p}(z))U_{p}]\xrightarrow[n\rightarrow+\infty]{}0\right]$
$\geq$ $\mathbb{P}\left[\sum_{p\geq 1}\left|\frac{1}{p}\operatorname{Tr}[(Q_{p}(z)-R_{p}(z))U_{p}]\right|^{2+\varepsilon/2}<+\infty\right]$
$=$ $1$
@@
Then, Lemma A.1 needs to be combined with the following one in order to prove Theorem 3.1.
@@
Let us define $\mathbf{T}(z)$ and $\mathbf{R}(z)$ as in Theorem 3.1 and Lemma A.1. Then one has that $\mathbf{R}(z)\sim\mathbf{T}(z),$ for all $z\in\mathcal{D}=\{z\in\mathbb{C}^{+},\frac{|z|}{|\Im(z)|}<2,|\Im(z)|>4d\gamma_{max}^{2}\}$.
@@
Since $T_{p}(z)$ and $R_{p}(z)$ are diagonal matrices, the following equations hold true $\mathrm{Tr}[(R_{p}(z)-T_{p}(z))U_{p}]=\sum_{i=1}^{p}(R_{p}^{(i)}(z)-T_{p}^{(i)}(z))U_{p}^{(i)}$, where $(U_{p})_{p\geq 1}$ denotes a family of deterministic matrices satisfying (3.3), and
@@
We know from [HLN07][Proposition 5.1], that $|R_{p}^{(i)}(z)T_{p}^{(i)}(z)|\leq\frac{1}{|\Im(z)|^{2}}$.
Moreover [HLN07][Equation (6.15)] states that
@@
Therefore, we obtain from these inequalities and (3.3) that
$\mathbb{E}\left[\sum_{p\geq 1}\left|\frac{1}{p}\operatorname{Tr}[(T_{p}(z)-R_{p}(z))U_{p}]\right|^{2+\delta/2}\right]$
$\leq$ $\sum_{p\geq 1}\frac{1}{p}\sum_{i=1}^{p}|U_{p}^{(i)}|^{2+\delta/2}\mathbb{E}\left[|T_{p}^{(i)}(z)-R_{p}^{(i)}(z)|^{2+\delta/2}\right]$
$\leq$ $\sum_{p\geq 1}\frac{K^{2+\delta/2}K_{2}}{|\Im(z)|^{2}p^{1+\delta/4}}\leq\sum_{p\geq 1}\frac{\widetilde{K}}{p^{1+\delta/4}}<+\infty,$
@@
Proof of Theorem 3.1.
[…]
@@
Hence, for each compact subset $C\subset\mathbb{C}\setminus\mathbb{R}^{+}$, $f_{p}$ is uniformly bounded on $C$, that is $|f_{p}(z)|\leq\frac{2K}{\delta_{C}}$, where $\delta_{C}$ is the distance between $C$ and $\mathbb{R}^{+}$. Then, by the normal family Theorem [Rud87][Theorem 14.6] there exists a sub-sequence $f_{p_{k}}$ which uniformly converges to $f^{\ast}$ that is an analytical function on $\mathbb{C}\setminus\mathbb{R}^{+}$. Let $(z_{k})_{k\in\mathbb{N}}$ be a sequence with an accumulation point in $\mathcal{D}$. Then, for each $k$, $f_{p}(z_{k})\rightarrow 0$ with probability one as $p$ tends to +$\infty$+ ∞. This implies that $f^{\ast}(z_{k})=0$ for each $k$. We finally obtain that $f^{\ast}$ is identically zero on $\mathbb{C}\setminus\mathbb{R}^{+}$. Therefore $f_{p}$ converges uniformly to zero which proves that $\frac{1}{p}\operatorname{Tr}[Q_{p}(z)U_{p}]\sim\frac{1}{p}\operatorname{Tr}[T_{p}(z)U_{p}]$ for $z\in\mathbb{C}\setminus\mathbb{R}^{+}$.
[…]
@@
Since we have proved that $\frac{1}{p}\operatorname{Tr}[Q_{p}(z)U_{p}]\sim\frac{1}{p}\operatorname{Tr}[T_{p}(z)U_{p}]$ holds on $\mathbb{C}\setminus\mathbb{R}^{+}$, we obtain that for all $w\in\rho$ that $\lim_{n\to\infty,\;p/n\to c}\frac{1}{p}\mathrm{Tr}[(Q_{p}(w)-T_{p}(w))U_{p}]=0.$ Moreover, for $w\in\mathbb{C}\setminus\mathbb{R}^{+}$, we have from (A.2) that
@@
Since $\frac{2K}{\mathrm{dist}(\rho,\mathbb{R}^{+})|w-z|^{2}}$ is integrable on $\rho$, using the dominated convergence theorem, we obtain that $\frac{1}{p}\operatorname{Tr}[Q_{p}^{\prime}(z)U_{p}]\sim\frac{1}{p}\operatorname{Tr}[T^{\prime}(z)U_{p}]$ holds for $z\in\mathbb{C}\setminus\mathbb{R}^{+}$ which completes the proof of Theorem 3.1. ∎